USE MATLAB TO SOLVE THIS PROBLEM

USE MATLAB TO SOLVE THIS PROBLEM

USE MATLAB TO SOLVE THIS PROBLEM

USE MATLAB TO SOLVE THIS PROBLEM

3. The range, r, of a cannon projectile under the influence of gravity, g-9.81 m/s, and fired with muzzle velocity vo 10 m/s is determined by the angle its barrel makes with the ground, t. From your Physics classes you will recall that the range is calculated as: r() = 2(vo)2/g sin t cos() . If a wind of velocity 2 m/s2/ blows against the project, what will be the r(0)? Assuming that 9 is a random variable uniformly distributed between O and T/2.
Answer #1

here resultant velocity v=v0-wind velocity=10-2=8 m/s

angle uniformly distributed between 0 and 90 deg that can be taken as0+90/2=45 deg/9=9.81m/s2

range of aprojectile r=v2sin(2\theta)/g

r=82sin(2*45)/9.8=6.523m